# Stratification governs the Existence of Surface-Intensified Eastward Jets in Turbulent Gyres without Bottom Friction

### v1

@@ 6 The Zonostrophic Regime ($10\lesssim\xi$): Freely Decaying Turbulence on Eastward Flow
We argue that this stretching is responsible for the apparition of multiple eastward jets and the creation of potential vorticity staircases in our simulations at $\xi=10$, $33$ and $100$. The scaling in equation (14) is confirmed in figure 8C. Following the idea of non-local generation of turbulence, $U_{turb}$ was calculated close to the western boundary and the jet scale and $U_{m}$ in the bulk (figure 8A, ”turbulence” and ”staircases” respectively). To further support the enhanced stretching on an imposed background flow, the scarce data available from the basin-scale runs was complemented with periodic simulations of decaying surface-intensified turbulence with an imposed background flow. Although slightly offset, they also follow the jet spacing given by equation (14). For further details on the periodic runs, extraction of $U_{turb}$, $U_{m}$ and the jet scale the reader may consult Appendix D.
@@ 7 Discussion and Conclusion
In the 1.5-layer model, the eastward jet emerges similarly to the two-layer case when $\xi$ increases above one. However, this process also leads to westward jets, which are not present in the two-layer model. We showed that these westward jets are disintegrated by baroclinic instability, while eastward jets remain baroclinically stable, explaining the establishment of the western boundary current extension as a single eastward jet.
In this study we investigated values of $\delta$ smaller than the usual oceanic configuration, but the flow dynamics at small $\delta$ remain relevant for understanding the classical configuration. Despite the asymptotically small value of $\delta$, the flow of the reference case shares many properties with the flow obtained with $\delta\sim 0.2$, free-slip boundary conditions and bottom friction (Holland, 1978) (see also appendix A). The underlying dynamical similarity between the classical flow configuration and the regime at $\delta=0.01$ becomes apparent in the results of linear stability analysis. Both configurations feature unstable westward Sverdrup flow, but eastward Sverdrup flow and inertial boundary currents are stable.
A common concern with quasi-geostrophic models of ocean gyres is how sensitive they are to boundary conditions. While free-slip boundary conditions tend to produce more intense eastward jets (Haidvogel et al., 1992; Deremble et al., 2011; Nasser et al., 2023), we showed that jets still form in a no-slip configuration if the stratification is appropriately chosen. One reason for this insensitivity to boundary conditions observed here might be the stability of the inertial western boundary layer in the two-layer system (Fig. 5), which is easier to model with free-slip bounary conditions (Ierley and Young, 1991).
@@
We conclude that the minimal ingredients for the emergence of a coherent eastward jet extending from western boundary layers in turbulent quasi-geostrophic gyres are: (i) a criticality parameter $\xi$ that is sufficiently large to permit baroclinic instability but sufficiently small to leave surface-intensified Sverdrup flow intact, (ii) a layer depth aspect ratio $\delta$ sufficiently small for baroclinic instability to permit eastward jets, but release energy towards eddies in adjacent areas of westward recirculation, and (iii) an efficient energy sink which suppresses unrealistic eddy-driven flows.
@@ Acknowledgements.
$\beta\frac{\partial\psi_{bt}}{\partial x}=\frac{\nabla\times\tau}{H_{1}+H_{2}}$ (15)
$\beta\frac{\partial\psi_{bc}}{\partial x}-\frac{1}{L_{d}^{2}}J(\psi_{bt},\psi_{bc})=\frac{\nabla\times\tau}{H_{1}}+\nu\frac{\nabla^{2}\psi_{bc}}{L_{d}^{2}}$ (16)
@@
In order to compare jet spacing in the zonostrophic regime with $L_{eff}^{Rh}$, we measured $U_{turb}$ as the average fluctuation speed along the meridional line $L/4<y<3L/4$, $x=3\delta_{I}$ and $U_{m}$ was measured as the mean transport across the line $L/4<y<3L/4$, $x=L/2$ divided by $L/2$ (figure 8A, lines entitled ”turbulence” and ”staircases” respectively). In order to determine the number of jets, we calculate the histogram of the potential vorticity along the Hovmöller diagram at mid-basin (figure 8B). Finally, to obtain the jet spacing, the meridional size over which the histogram was obtained is then divided by the number of distinct peaks that exceed the standard deviation of the histogram, corresponding to the number of distinct homogenized regions.
@@
$\psi_{1}(t=0)$ $=\psi^{\prime}_{1}-U_{m}y$ (28)
$\psi_{2}(t=0)$ $=0$ (29)
$q_{1}(t=0)=q^{\prime}_{1}+\frac{1-\delta}{L_{d}^{2}}U_{m}y$ (30)
$q_{2}(t=0)=-\frac{\delta}{L_{d}^{2}}U_{m}y$ (31)

### v2

@@ 6 The Zonostrophic Regime ($10\lesssim\xi$): Freely Decaying Turbulence on Eastward Flow
We argue that this stretching is responsible for the appearance of multiple eastward jets and the creation of potential vorticity staircases in our simulations at $\xi=10$, $33$ and $100$. The scaling in equation (16) is confirmed in figure 8C. Following the idea of non-local generation of turbulence, $U_{turb}$ was calculated close to the western boundary and the jet scale and $U_{m}$ in the bulk (figure 8A, “turbulence” and “staircases” respectively). To further support the enhanced stretching on an imposed background flow, the scarce data available from the basin-scale runs was complemented with periodic simulations of decaying surface-intensified turbulence with an imposed background flow. Although slightly offset, they also follow the jet spacing given by equation (16). For further details on the periodic runs, extraction of $U_{turb}$, $U_{m}$ and the jet scale the reader may consult Appendix D.
@@ 7 Discussion and Conclusion
In the 1.5-layer model, the eastward jet emerges similarly to the two-layer case when $\xi$ increases above one. However, this process also leads to westward jets, which are not present in the two-layer model. We showed that these westward jets are disintegrated by baroclinic instability, explaining the establishment of the western boundary current extension as a single eastward jet.
In this study we investigated values of $\delta$ smaller than the usual oceanic configuration, but the flow dynamics at small $\delta$ remain relevant for understanding the classical configuration. Despite the asymptotically small value of $\delta$, the flow of the reference case shares many properties with the flow obtained with $\delta\sim 0.2$, free-slip boundary conditions and bottom friction (Holland, 1978) (see also appendix A). The underlying dynamical similarity between the classical flow configuration and the regime at $\delta=0.01$ becomes apparent in the results of linear stability analysis. Both configurations feature unstable westward Sverdrup flow, but eastward Sverdrup flow and inertial boundary currents are stable. Less extreme surface-intensification will likely shift the exact location of the eastward jet in the parameter space of a more realistic model. If understood as the transitory state between a vortex gas and a zonostrophic regime (figure 4), its formation may be altered for example by bottom friction (Held and Larichev, 1996; Gallet and Ferrari, 2021) or topographic stretching close to the western boundary (Boland et al., 2012).
A common concern with quasi-geostrophic models of ocean gyres is how sensitive they are to boundary conditions. While free-slip boundary conditions tend to produce more intense eastward jets (Haidvogel et al., 1992; Deremble et al., 2011; Nasser et al., 2023), we showed that jets still form in a no-slip configuration if the stratification is appropriately chosen. One reason for this insensitivity to boundary conditions observed here might be the stability of the inertial western boundary layer in the two-layer system (Fig. 5), which is easier to model with free-slip boundary conditions (Ierley and Young, 1991).
@@
We conclude that the minimal ingredients for the emergence of a coherent eastward jet extending from western boundary layers in turbulent quasi-geostrophic gyres are: (i) a criticality parameter $\xi$ that is sufficiently large to permit baroclinic instability ($\xi>1$) but sufficiently small to leave surface-intensified Sverdrup flow intact ($\xi<1/\delta$), (ii) a layer depth aspect ratio $\delta$ sufficiently small for baroclinic instability to exhibit zonal asymmetry but release energy towards eddies in adjacent areas of westward recirculation, and (iii) an efficient energy sink which suppresses unrealistic eddy-driven flows.
A natural extension of the present work is to use these findings to understand the dynamical regimes observed in the world ocean. The first step will be to go beyond the quasi-geostrophic model, which assumes a prescribed, horizontally homogeneous stratification profile. The second step will be to generalize the results obtained from a two-layer model to continuously stratified models. We are currently investigating these two issues using both climatological data and primitive equations model runs.
@@ Acknowledgements.
Figure 10: Snapshots of upper layer (A, B) and lower layer (C, D) potential vorticity of the free-slip, bottom friction configuration as in figure 9 but in the strong $\left(\xi=1.3\times 10^{-2}\right)$ and weak $\left(\xi=1.3\times 10^{2}\right)$ stratification regime. In both regimes the jet disappears and is replaced by energetic vortices close to the western boundary.
The jet also disappears in this configuration when entering the regimes of strong and weak stratification regimes in this configuration, as illustrated in figure 10. Strong vortices appear close to the western boundary current as in the no-slip configuration where energy is dissipated by vorticitiy filaments. In the strong stratification regime the flow is completely confined to the upper layer, while in the weak stratification regime the flow becomes completely barotropic. The flow dynamics of these regimes are investigated in Greatbatch and Nadiga (2000).
@@
$\beta\frac{\partial\psi_{bt}}{\partial x}$ =×$=\frac{\nabla\times\tau}{H_{1}+H_{2}}$ (17)
$\beta\frac{\partial\psi_{bc}}{\partial x}$ $=\frac{1}{L_{d}^{2}}J(\psi_{bt},\psi_{bc})+\frac{\nabla\times\tau}{H_{1}}+\nu\frac{\nabla^{2}\psi_{bc}}{L_{d}^{2}}$ (18)
@@
In order to compare jet spacing in the zonostrophic regime with $L_{eff}^{Rh}$, we measured $U_{turb}$ as the average fluctuation speed along the meridional line $L/4<y<3L/4$, $x=3\delta_{I}$ and $U_{m}$ was measured as the mean transport across the line $L/4<y<3L/4$, $x=L/2$ divided by $L/2$ (figure 8A, lines entitled “turbulence” and “staircases” respectively). In order to determine the number of jets, we calculate the histogram of the potential vorticity along the Hovmöller diagram at mid-basin (figure 8B). Finally, to obtain the jet spacing, the meridional size over which the histogram was obtained is then divided by the number of distinct peaks that exceed the standard deviation of the histogram, corresponding to the number of distinct homogenized regions.
@@
$\psi_{1}(t=0)=\psi^{\prime}_{1}-U_{m}y$ $\ q_{1}(t=0)=q^{\prime}_{1}+\frac{1-\delta}{L_{d}^{2}}U_{m}y$ (30)
$\psi_{2}(t=0)=0$ $\ q_{2}(t=0)=-\frac{\delta}{L_{d}^{2}}U_{m}y$ (31)